OpenStudy (anonymous):

Will Medal Which of the following is equal to the square root of the cube root of 5? 5 to the power of 1 over 3 5 to the power of 1 over 6 5 to the power of 2 over 3 5 to the power of 3 over 2

OpenStudy (anonymous):

@taiyanna.nicole @AbdullahM @AutumnRoseT

OpenStudy (abdullahm):

Hi

OpenStudy (anonymous):

hi i need help quick

OpenStudy (abdullahm):

Hint: \(\sf\Huge \sqrt[x]{y}=y^{\frac{1}{x}}\)

OpenStudy (anonymous):

so a

OpenStudy (abdullahm):

Your question is: \(\sf\Huge \sqrt[2]{\sqrt[3]{5}}=?\)

OpenStudy (anonymous):

no its square root thing then a 3 then another squareroot thing inside the other with the 3 on top of it with a 5 inside that one

OpenStudy (anonymous):

so yes without the 2

OpenStudy (abdullahm):

Could you take a screenshot of it?

OpenStudy (anonymous):

its the same thing without the 2

OpenStudy (abdullahm):

Like this? \(\sf\Huge\sqrt[3]{5}=?\)

OpenStudy (abdullahm):

@DaBest21

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

just add a square root thing on top of the 3

OpenStudy (anonymous):

\[\sqrt{3\sqrt{5}}\]

OpenStudy (anonymous):

@AbdullahM r u there

OpenStudy (abdullahm):

Btw, \(\sf\Huge\sqrt{x} = \sqrt[2]{x}\) There is no difference.

OpenStudy (anonymous):

ok srry

OpenStudy (anonymous):

soncan u help me plz

OpenStudy (abdullahm):

\(\color{blue}{\text{Originally Posted by}}\) @AbdullahM Your question is: \(\sf\Huge \sqrt[2]{\sqrt[3]{5}}=?\) \(\color{blue}{\text{End of Quote}}\) What is \(\sf\Large\sqrt[3]{5}\) using my hint above.

OpenStudy (anonymous):

1.709

OpenStudy (abdullahm):

\(\color{blue}{\text{Originally Posted by}}\) @AbdullahM Hint: \(\sf\Huge \sqrt[x]{y}=y^{\frac{1}{x}}\) \(\color{blue}{\text{End of Quote}}\) No, we don't want it in decimal. Use this hint ^^^

OpenStudy (anonymous):

5 1/3

OpenStudy (anonymous):

@AbdullahM help

OpenStudy (anonymous):

@brigirl1212 @Goku-Kai

OpenStudy (abdullahm):

correct

OpenStudy (abdullahm):

now we have: \(\sf\Huge \sqrt[2]{5^{\frac{1}{3}}}\)

OpenStudy (abdullahm):

change the remaining squareroot and what do we get?

OpenStudy (anonymous):

@AbdullahM help

OpenStudy (anonymous):

5 1/3 1/2

OpenStudy (anonymous):

5 1/3 1/2

OpenStudy (abdullahm):

great

OpenStudy (abdullahm):

so now we have \(\sf\Large (5^{\frac{1}{3}})^{\frac{1}{2}}=?\)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

now what @AbdullahM

OpenStudy (anonymous):

@AbdullahM

OpenStudy (anonymous):

@AbdullahM

OpenStudy (anonymous):

@AbdullahM

OpenStudy (anonymous):

now what

OpenStudy (abdullahm):

Hint: \(\sf\Large (a^b)^c = a^{b\times c}\)

OpenStudy (anonymous):

5 1/6 ?

OpenStudy (abdullahm):

correct and that's your answer

OpenStudy (anonymous):

thanks

OpenStudy (abdullahm):

np :)